# Debunking the CUDA Myth Towards GPU-based AI Systems : Evaluation of the Performance and Programmability of Intel’s Gaudi NPU for AI Model Serving

### v1

@@ 4.2. Performance Optimization at the High-level PyTorch: A vLLM Case Study
Figure 16 shows vLLMopt’s effect in improving PagedAttention and end-to-end LLM performance. On average, vLLMopt achieves 7.4 improvement in PagedAttention throughput over vLLMbase when the fraction of zero-padded indices is 0% (Figure 16(a)). This result highlights the efficacy of vLLMopt’s PyTorch-level performance optimizations and graph compiler’s effectiveness in pipelining PagedAttention’s operations across MME and TPC. Furthermore, the experiment in Figure 16(b) shows that PagedAttention’s throughput improves by up to 55.7 (average 21) as the fraction of zero-padded indices increases, emphasizing the importance of eliminating redundant operations. Despite these improvements, vLLMopt still falls short of A100, achieving an average of 45% of A100’s PagedAttention throughput. However, due to Amdahl’s law and Gaudi-2’s performance gains in GEMM operations (Section 3.2), the vLLMopt-based Gaudi-2 demonstrates comparable end-to-end performance and TPOT to A100, as shown in Figure 16(c).
@@ 5. Discussion and Future Work
(Future Work) Because NVIDIA GPUs are the de facto standard in AI systems, we focused on comparing Gaudi against NVIDIA GPUs, leaving out the comparison with AMD GPUs (AMD, 2024) or other NPUs (Jouppi et al., 2023; Amazon Web Services, 2024a). Possible future work includes evaluating Gaudi against these alternative platforms. Additionally, Intel claims that Gaudi NPUs are competitive to NVIDIA GPUs for training large-scale AI models requiring hundreds to thousands of devices. Analyzing Gaudi’s competitive edge against NVIDIA GPUs in training scenarios is part of our immediate future work.

### v2

@@ 4.2. Performance Optimization at the High-level PyTorch: A vLLM Case Study
Figure 17 shows vLLMopt’s effect in improving PagedAttention and end-to-end LLM performance. On average, vLLMopt achieves 7.4 improvement in PagedAttention throughput over vLLMbase when the fraction of zero-padded indices is 0% (Figure 17(a)). This result highlights the efficacy of vLLMopt’s PyTorch-level performance optimizations and graph compiler’s effectiveness in pipelining PagedAttention’s operations across MME and TPC. Furthermore, the experiment in Figure 17(b) shows that PagedAttention’s throughput improves by up to 55.7 (average 21) as the fraction of zero-padded indices increases, emphasizing the importance of eliminating redundant operations. Despite these improvements, vLLMopt still falls short of A100, achieving an average of 45% of A100’s PagedAttention throughput (Figure 17(c)). However, due to Amdahl’s law and Gaudi-2’s performance gains in GEMM operations in MLP layers (Section 3.2), the vLLMopt-based Gaudi-2 demonstrates similar end-to-end performance (Figure 17(d)) to A100, with comparable sensitivity to SLO (service level objective) oriented metrics when sweeping the inference server’s maximum batch size (i.e., the change in TTFT (Time-To-First-Token) vs. TPOT (Time-Per-Output-Token) (Figure 17(e))).
@@ 5. Discussion and Future Work
(Future Work) Because NVIDIA GPUs are the de facto standard in AI systems, we focused on comparing Gaudi against NVIDIA GPUs, leaving out the comparison with AMD GPUs (AMD, 2024) or other NPUs (Jouppi et al., 2023; Amazon Web Services, 2024a). Possible future work includes evaluating Gaudi against these alternative platforms. Additionally, Intel claims that Gaudi NPUs are competitive to NVIDIA GPUs for training large-scale AI models requiring hundreds to thousands of devices. Analyzing Gaudi’s competitive edge against NVIDIA GPUs in training scenarios is part of our immediate future work. Furthermore, AMD’s recently announced Strix Halo (Ryzen AI Max) processor (AMD, 2025) integrates Zen CPU cores, RDNA GPU architecture, and XDNA 2 NPU into a single SoC, offering a distinct alternative to both NVIDIA GPUs and Intel Gaudi NPUs. Unlike Gaudi, which is designed primarily for large-scale distributed AI workloads, Strix Halo targets efficient on-device AI inference with its XDNA 2 NPU, while also providing a powerful integrated GPU for mixed AI and graphics workloads. Future work could explore how Strix Halo’s unified memory architecture and heterogeneous compute capabilities compare to Gaudi NPUs and NVIDIA GPUs for AI serving.